Find all solutions to the equation 3x^(1/6)+2x^(1/3)-2=0
Answer by stanbon(75887)

Let w=x^(1/6) the w^2=x^(1/3)
Rewrite the equation using "w" as the variable, as follows:
2w^2+3w-2=0
(2w-1)(w+2)=0
w=1/2 or w=-2
Now, switch back to the "x" variable, as follows:
x^(1/6)=1/2 so x=(1/2)^6 = 1/64
or
x^(1/6)=-2 so x=(-2)^6= 64
Cheers,
Stan H.